OpenStudy (anonymous):

15. Simplify 6P2 A. 30 B. 15 C. 12 D. 720

OpenStudy (anonymous):

\[ ^{\color{red}{6}}P_{\color{blue}{2}} = \frac{\color{red}{6}!}{(\color{red}{6}-\color{blue}{2})!} = \frac{6!}{4!} = \frac{6\times 5 \times \cancel{4!}}{\cancel{4!}} = 6\times 5 \]

OpenStudy (anonymous):

:D Thanks!